Ticket Number: HEA-NOC/20060825-2 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20060825 09:01 UTC+1 Problem Start: 20060825 02:28 UTC+1
Ticket Update: 20060825 10:07 UTC+1 Problem End:
Site/line: Crumlin VEC, HEA, IRCSET, Kilester VEC, Liberties VEC,
Monitoring tools detect loss of connectivity to the HEA, IRCSET and CDVECs
since 2:30 AM. NBE euipment at the same site is also unreachable.
No traffic to or from affected sites.
20060825 08:56 AH Suspected onsite problem. Contacted site contacts.
20060825 10:04 AH Event is dependent ona wider Colt outage, some of which
is also detailed in ticket HEA-NOC/20060825-2.
This ticket can be monitored at http://www.hea.net/tickets/20060825-2